OpenStudy (anonymous):

solve by substitution 0.4x=1 3x-2y=9 how do i solve this problem im having issues trying to solve it please help!

OpenStudy (mrdoe):

so plug x=2.5 into your seconf eq instead of x and solve for y

OpenStudy (mrdoe):

second*

OpenStudy (anonymous):

1/0.4=2.5 x=2.5 3(2.5)-2y=9 7.5-2y=9 -2y=1.5 -2y/2=1.5/2 y=-0.5

OpenStudy (anonymous):

but what about the x

OpenStudy (mrdoe):

you already have x by itself in the forst eq: x=1/0.4=2.5

OpenStudy (mrdoe):

first*

OpenStudy (anonymous):

so if i put it in pairs it would be (2.5,-0.75)

OpenStudy (anonymous):

right?